# Long-Horizon Planning for Multi-Agent Robots in Partially Observable Environments

### v1

@@ 7 Limitations and Future Work
Limited spatial reasoning: Although we use both textual descriptions and visual features to guide the language model’s actions, it still lacks the ability to reason about the spatial features of the environment. Spatial reasoning is crucial in scenarios such as navigating around obstacles to reach an object, or determining the shortest path to collect multiple items scattered across different locations. One way to address this limitation is to inject information about the 3D world into the LM, as done in (Hong et al., 2023b), which is an interesting direction for future work.
@@ Acknowledgements
We would like to thank Keerthana Gopalakrishnan, Sydney Dolan, Jasmine Aloor, and Victor Qin for helpful discussions and feedback. OpenAI credits for GPT-4 access was provided through OpenAI’s Researcher Access Program. The research was sponsored by the United States Air Force Research Laboratory and the Department of the Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ B.2 Action Space
Explore action $\mathcal{A}_{EXP}$: The explore action is carried out by the heuristic mentioned in Algorithm 1. We use the clip-vit-large-patch14-336 model for the CLIP weights which we download from https://huggingface.co/openai/clip-vit-large-patch14-336.
@@
Algorithm 1 Exploration Heuristic
@@
Appendix C Task Types
@@
  • Implicit target and object types: Here both the object type and the target are explicitly defined. E.g. Clear the floor by placing the items at their appropriate positions. Here the model is expected to keep items like pens, book, laptop on the study table, litter in the trash can, etc.
@@
Appendix D Pseudocode for LLaMAR
@@
Appendix E Baselines
@@
While there are a lot of impressive LLM-based multi-agent planners as mentioned in Table 1, they vary in the assumptions about access to information about the environment. We were not able to find the official codebase for the Safe Multi-Agent Planning with Conformal Prediction (Wang et al., 2024) and TwoStep (Singh et al., 2024). We describe the prompts used for our model as well as every baseline. Note that we show the prompt for the 2-agent case, but it is easily modified to generalize to the $n$-agent case. The italics and bolding added for emphasis.
E.1 LLaMAR
@@
description of what failed,, Bob’s observation: list of objects the Bob is observing,
@@ E.6 CoELA
Appendix F Open Source VLMs
@@
  • Idefics 2 (Laurençon et al., 2023, 2024): We use the 8B base model fine-tuned on a mixture of supervised and instruction datasets (text-only and multimodal datasets) from HuggingFace. The weights were downloaded from https://huggingface.co/HuggingFaceM4/idefics2-8b with the commit #2c031da2dc71f3ac989f9efa9b8ff476df3842c0. We chose Idefics because it is able to take multiple images as input similar to GPT-4V and reason on them.
  • LLaVA (Liu et al., 2023b): We use the 7B model t trained by fine-tuning LLaMA/Vicuna on GPT-generated multimodal instruction-following data. The weights were downloaded from https://huggingface.co/llava-hf/llava-1.5-7b-hf with the commit # 05ae2434cbb430be33edcba0c5203e7023f785b7.
  • CogVLM (Hong et al., 2023a): We use the 18B model. The weights were downloaded from https://huggingface.co/THUDM/cogagent-chat-hf with the commit # d519da3b191401234f4bd86ce1c287c61bc276a3.
Appendix G SentenceBERT fine-tuning

### v2

@@ 7 Limitations and Future Work
Limited spatial reasoning: Although we use both textual descriptions and visual features to guide the language model’s actions, it still lacks the ability to reason about the spatial features of the environment. Spatial reasoning is crucial in scenarios such as navigating around obstacles to reach an object, or determining the shortest path to collect multiple items scattered across different locations. One way to address this limitation is to inject information about the 3D world into the LM, as done in Hong et al. (2023b), which is an interesting direction for future work.
@@ Acknowledgements
We would like to thank Keerthana Gopalakrishnan, Sydney Dolan, Jasmine Aloor, and Victor Qin for helpful discussions and feedback. OpenAI credits for GPT-4 access was provided through OpenAI’s Researcher Access Program. The research was sponsored by the Department of the Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ B.2 Action Space
Explore action $\mathcal{A}_{EXP}$: The explore action is carried out by the heuristic explained in Algorithm 1 and Figure 4. We use the clip-vit-large-patch14-336 model for the CLIP weights which we download from https://huggingface.co/openai/clip-vit-large-patch14-336.
The action space consists of several executable actions, including Move(<direction>), Rotate(<direction>), LookUp(<angle>), LookDown(<angle>), NavigateTo($<$object_id$>$), Pickup($<$object_id$>$), Put($<$receptacle_id$>$), and others. These actions can be combined in numerous ways, as the $<$object_id$>$, <directions>, and <angles> can vary significantly, resulting in a combinatorially large action space. To address this complexity, we do not impose constraints on the language model (LM) modules to select actions in a predefined format from this set. Instead, we allow for free-form language outputs, offering greater expressivity. The SentenceBERT module then maps these free-form natural language outputs to executable actions within the environment.
@@
Algorithm 1 AI2Thor Exploration Heuristic
@@
Appendix C MAP-THOR Task Types
@@
  • Implicit target and object types: Here both the object type and the target are implicitly defined. E.g. Clear the floor by placing the items at their appropriate positions. Here the model is expected to keep items like pens, book, laptop on the study table, litter in the trash can, etc.
@@
Appendix D Search & Rescue Environment (SAR)
The Search & Rescue environment consists of multiple agents in an unknown environment that has multiple wildfires and missing personnel in the environment. The agents are tasked to extinguish all the fires before they spread and rescue the missing humans. Here, each fire is composed of a large flammable region with a fixed set of sources that spread through time. The higher the intensity, the faster the fire will spread. The fires can be of class A or B, extinguished through the use of water and sand respectively. Both these resources can be collected through resource reservoirs spread geographically. Each person is initially stranded in an unknown location. The goal is to rescue and transport each person to a drop-off location (known apriori). The person must have two agents simultaneously carrying them to be transported to the drop-off location.
Figure 5: The search & rescue environment consists of multiple drones in an unknown environment with missing people, fires of different types, and water and sand reservoirs.
We provide a justification to why this task is relevant to long-term planning and multi-agent collaboration below.
D.1 Multi-agent Nature
  • Scalability - Unlike the AI2THOR environment, the search & rescue environment is more spacious and hence prevents congestion amongst agents. With the addition of multiple parallel emergency disaster scenarios at once, the environment scales comfortably to more agents. Additionally, the complexity slowly increases as the model has to coordinate between more actors and competing emergencies.
  • Tasks requires explicit cooperation -
    • To move a group of humans, at least 2 agents are required. Thus, explicit multi-agent collaboration is required.
    • Due to the time pressure, in order to successfully stop the fire, agents must all effectively collaborate in collecting and strategically using resources in high-intensity regions.
  • Exploration & Task Assignment - There is an explicit tradeoff between allocating agents towards exploring to find lost people and fighting the current fires.
D.2 Long-term Planning
  • Dependency Chain: To resolve a fire it requires that the source is identified, the type of fire is identified, and the appropriate resources to stop it are acquired and then used.
  • Uncertainty: Inherent uncertainty as to where the lost people can be found, and at what point in the timeline they will.
  • Irreversibility and forced consequences - With a fire that spreads, present actions have irreversible future consequences. A balance needs to be struck between fighting the fire’s source (to stop it from continuing) versus a periphery (to prevent geographic spread).
D.3 Description of Scenes
We evaluate LLaMAR in five different scenarios in the search & rescue environment. Here each scene is evaluated on 5 different random seeds.
  • Scene 1: This consists of 1 Type A fire with 2 initial sources, 1 Type B fire with 1 initial source, and 1 lost person at a random location in the environment.
  • Scene 2: This consists of 1 Type A fire with 1 initial source, 1 Type B fire with 2 initial sources, and 1 lost person at a random location in the environment.
  • Scene 3: This consists of 2 Type A fires each with 1 initial source, 1 Type B fire with 1 initial source, and 1 lost person at a random location in the environment.
  • Scene 4: This consists of 1 Type A fire with 3 initial sources, and 1 lost person at a random location in the environment.
  • Scene 5: This consists of 1 Type A fire with 1 initial source, 1 Type B fire with 1 initial source, and 2 lost persons at random locations in the environment.
D.4 Observation Space
The agent’s observation $\mathcal{O}=\mathcal{O}_{G}\cup\mathcal{O}_{L}\cup\mathcal{O}_{L}$ is a union of the following observations.
Global Observations $\mathcal{O}_{G}$ consists of all the objects that are globally visible by the agent:
  • Fires: If visible, fire name, type, and average intensity.
  • Fire Regions: If fire visible, fire region name, type, and average intensity.
  • Reservoir: If visible, reservoir name, and reservoir resource type.
  • Deposit: If visible, deposit name, inventory of all the resources (water, sand) and persons in deposit.
  • Person: If visible, person name, carried or not status, dropped-off or not status.
  • Agent inventory: List of all the resources (water, sand) and person (being carried) in the agent’s inventory.
Local Observations $\mathcal{O}_{L}$ consists of all the objects that are visible in grid cells adjacent to the agent:
  • For all the directions <direction> in (Up, Down, Left, Right, and Center), output of the following.
  • <direction>: At direction <direction>, either ‘Empty’ if there is no object, ‘Flammable’ along with corresponding intensity & fire name if object is a part of a fire, or ‘Obstacle’ for any other object.
Names $\mathcal{O}_{N}$ consists of a list of the name of all visible and interactable objects.
D.5 Action Space
The action space $\mathcal{A}$ consists of navigation actions $\mathcal{A}_{NAV}$, interaction actions $\mathcal{A}_{INT}$, exploration action $\mathcal{A}_{EXP}$.
  • Move(<direction>): Moves the agent to the neighboring grid cell in the specified direction where <direction> can be one of (Up, Down, Left, Right, and Center)
  • NavigateTo(<targetID>): Moves the agent next to the location of the object <targetID> if <targetID> is visible.
  • Carry(<person>): Makes agent carry a person <person> if <person> is visible and interactable. The person is successfully ‘group’ carried, if at least the required number of agents successfully does the ‘Carry(<person>)’ action. If carry action is successful, all other resources in agent’s inventory are dropped.
  • DropOff(<person>, <deposit>): Drops off person <person> at location <deposit>. This action is only successful if the person has been ‘group’ carried, all the agents carrying the person have the deposit be visible and interactable, and after all the agents do this action.
  • StoreSupply(<deposit>): Stores all the resources from the agent’s current inventory in the deposit <deposit>.
  • UseSupply(<fire>, <supply-type>): Uses all the supplies of type <supply-type> on the fire <fire> at the location the agent is at.
  • GetSupply(<deposit>, <supply-type>): Fills the remaining of the agent’s inventory space with the available supply of type <supply-type> in <deposit>.
  • GetSupply(<reservoir>): Collects 1 unit of supply from reservoir <reservoir> and stores it in the agent’s inventory.
Exploration action $\mathcal{A}_{EXP}$ consists of the following actions:
  • Explore(): Takes the agent in a direction of exploration as described by the heuristic exploration function described in algorithm 2.
Algorithm 2 SAR Exploration Heuristic
Input: Agent $A$, Environment $env$, Previous Direction $D$, Current Position $P$, Max steps $M$
Initialize: New direction $ND\leftarrow\emptyset$
1:$ND\leftarrow$ randomly choose an angle from $[0,2\pi)$ at $\pi/4$ intervals
2:while $ND=D$ or $ND\equiv D+\pi\mod{2\pi}$
3:          or a barrier exists at most $\frac{3}{4}M$ steps in direction $ND$ from $P$ do
4:     $ND\leftarrow$ randomly choose an angle from $[0,2\pi)$ at $\pi/4$ intervals
5:end while
6:Move agent $A$, $M$ steps in direction $ND$ in environment $env$
Appendix E Pseudocode for LLaMAR
@@
Appendix F Full Results
      Algorithm     LM     Success Transport Coverage Balance Steps
            Rate Rate
      Act     GPT-4V     0.33 (0.19, 0.49) 0.67 (0.59, 0.76) 0.91 (0.86,0.95) 0.59 (0.52, 0.66) 24.92 (22.12,27.73)
    ReAct     GPT-4V     0.34 (0.20, 0.49) 0.72 (0.63,0.80) 0.92 (0.86, 0.97) 0.67 (0.61, 0.73) 24.08 (21.27, 26.89)
    CoT     GPT-4V     0.14 (0.06, 0.28) 0.59 (0.51, 0.67) 0.87 (0.81, 0.92) 0.62 (0.56,0.69) 28.40 (26.91, 29.97)
    SmartLLM     GPT-4V     0.11 (0.05, 0.23) 0.23 (0.13, 0.31) 0.91 (0.80, 0.96) 0.45 (0.37, 0.52) 29.87 (26.20, 30.00)
    CoELA     GPT-4V     0.25 (0.10, 0.36) 0.46 (0.35, 0.56) 0.76 (0.67, 0.85) 0.73 (0.67, 0.80) 28.93 (27.77,30.00)
      LLaMAR     GPT-4     0.51 (0.36, 0.66) 0.85 (0.80, 0.91) 0.95 (0.91, 0.98) 0.83 (0.78, 0.86) 25.80 (23.72, 27.88)
    LLaMAR     LLaVA     0.54 (0.41, 0.65) 0.84 (0.71, 0.90) 0.91 (0.87, 0.98) 0.75 (0.64, 0.83) 26.21 (21.56, 28.97)
    LLaMAR     IDEFICS-2     0.57 (0.43, 0.67) 0.86 (0.74, 0.91) 0.94 (0.89, 0.98) 0.78 (0.65, 0.84) 25.27 (20.14, 28.37)
    LLaMAR     CogVLM     0.61 (0.47, 0.68) 0.89 (0.73, 0.95) 0.95 (0.89, 0.99) 0.80 (0.73, 0.86) 23.21 (20.57, 26.82)
    LLaMAR     GPT-4V     0.62 (0.46, 0.76) 0.87 (0.80, 0.93) 0.95 (0.91, 0.98) 0.82 (0.77, 0.87) 23.44 (20.88, 26.00)
    (w/o exploration)
    LLaMAR     GPT-4V     0.66 (0.50, 0.78) 0.91 (0.81,0.96) 0.97 (0.93,0.99) 0.82 (0.75,0.87) 21.87 (18.76,24.23)
    (w/ exploration)
Table 5: Comparison of evaluation metrics against baselines averaged across all tasks for the 2-agent MAP-THOR scenarios.
      Modules     Success Transport Coverage Balance Steps
    Used     Rate Rate
      Actor     0.33 0.67 0.91 0.59 24.92
        (0.19, 0.49) (0.59, 0.76) (0.86,0.95) (0.52, 0.66) (22.12,27.73)
    Planner +     0.45 0.78 0.92 0.69 24.87
    Verifier
    Planner +     0.67 0.91 0.97 0.84 22.81
    Corrector
    LLaMAR     0.66 0.91 0.97 0.82 21.87
        (0.50, 0.76) (0.81, 0.96) (0.93,0.99) (0.75, 0.87) (18.76, 26.43)
          MAP-THOR     SAR
    # of     Success Transport Coverage Balance Steps     Success Transport Coverage Balance Steps
    agents     Rate Rate     Rate Rate
      1     0.37 0.67 0.87 1.00 28.44     0.28 0.75 0.86 1.00 28
        (0.21, 0.51) (0.58, 0.74) (0.81, 0.90) (1.00,1.00) (25.23, 30.00)
    2     0.62 0.87 0.95 0.82 23.44     0.44 0.86 0.94 0.91 27.76
        (0.46, 0.76) (0.80, 0.93) (0.91, 0.98) (0.77,0.87) (20.88, 26.00)     (0.24, 0.65) (0.79, 0.94) (0.88, 0.99) (0.88, 0.95) (24.15, 30)
    3     0.70 0.91 0.98 0.66 21.30     0.68 0.92 0.96 0.80 21.88
        (0.55, 0.82) (0.85, 0.95) (0.95, 0.99) (0.61, 0.71) (18.60, 23.99)     (0.46, 0.85) (0.86, 0.98) (0.91,1.0) (0.73, 0.86) (17.83, 25.92)
    4     0.68 0.90 0.99 0.62 22.83     0.72 0.94 0.98 0.78 22.00
        (0.52, 0.79) (0.84, 0.94) (0.95, 0.99) (0.57, 0.68) (19.63, 25.69)     (0.50, 0.85) (0.88,0.98) (0.93, 1.00) (0.74, 0.83) (17.96, 26.03)
    5     0.62 0.90 0.99 0.54 22.91     0.74 0.96 1.00 0.73 24.52
        (0.46, 0.75) (0.85, 0.94) (0.97,1.00) (0.48, 0.59) (20.26, 25.57)     (0.52, 0.86) (0.94, 0.99) (1.0,1.0) (0.67, 0.79) (20.24,28.79)
Table 7: LLaMAR with more agents
Appendix G Failure Cases
One of the main motivations of our paper was to achieve better success rates in planning for a variety of tasks compared to previous approaches. While our method outperforms other baselines, we acknowledge that the success rate is still under the expectation for real-world deployment. We believe that our approach LLaMAR and the MAP-THOR benchmark would serve as a starting point for future research to improve the success rates in multi-agent embodied robotics tasks. Here we describe a few major types of failures in our experiments:
  • Mis-generalization: The agent can sometime fail to properly infer the level of abstraction to perform a task at, even if it should be clear from the action space. For example, for the washing tasks, the LM assumes that it must put the objects in the sink and add some, rather than just using the "CleanObject" action. We observe this error in the direction of performing actions more low-level than necessary, and not the other way around (more high-level than is feasible).
  • Mutual Interference (limited spatial reasoning): The agents sometimes block each other and thus fail to carry out actions such as placing an object on a receptacle, opening an object, etc. In particular, we see this behavior in the "put objects on the sofa" and the "put objects in a box" tasks, where the LM does not prevent the agents from blocking each other.
  • Improper Object Visibility (un-observability bias): The LM often fails to prioritize exploration for tasks that require objects not yet seen, simply due to it no being able to see it. This bias causes it to improperly assume that further exploration is not necessary, so it fails to find relevant objects.
  • SentenceBERT mismatch: The free-form natural language output from the Actor is incorrectly mapped to the executable action. Based on our experimental results, 96.7% of the time the free-form natural language was correctly mapped to the feasible actions. The 2 main reasons for failures in the few cases it failed were:
    • Incorrect mapping of objects: When the agents have not explored the environment enough and the Actor module suggests the agents interact with objects that are not yet available in the agents’ combined memory. Example: The free-form output “navigate to the table” was mapped to the action “NavigateTo(ArmChair)”.
    • Wrong object numeration: When there are multiple objects of the same type in the memory, the sentenceBERT maps “Cabinet_3" to “Cabinet_1" which sometimes leads to incorrect actions being executed for the completion of the plan.
  • Limited horizon: While each task is solvable by a single-agent human-controlled agent within our chosen episode length of $L=30$, this is under the assumption of no failures during control and maximum visibility. For the autonomous agents the visibility distance is restricted to 1.5 meters, so while a human could identify objects from afar (from the rendered images) the agents have to explore the environment under a stricter partial observability constraint.
    We hypothesize that a higher episode length would lead to higher metrics, but the cutoff was chosen for computational budget considerations. As can be observed in figure 6, the coverage and transportation rate metrics in particular seem to plateau around our chosen episode length $L$, so $L=30$ is appropriate.
Figure 6: Plots of the average transport rate (a),(d); coverage (b),(e); and success rate (c),(f) metrics vs. maximum horizon steps for the GPT-4V, LLaMAR algorithm on MAP-THOR. Shown for both the 2 agent (a),(b),(c); and 5 agent cases (d),(e),(f), along with an error region of one standard deviation.
G.1 Failure Cases
We have observed the following categories of failures for the LM on the SAR environment.
  • Incorrect Causal Ordering: The LM sometimes fails to execute the steps in the proper causal ordering. For instance, we have observed the agent trying to drop off a lost person before picking them up, or properly getting supplies to stop the fire but not navigating to the fire location before using them.
  • Sub-Optimal Action Sequence: Due to the time-constraints of the fire spread, it is not sufficient for the LM to do the sequence of tasks in order. They must also be done in a non-redundant, timely fashion. Otherwise, we have observed that the LM takes the right action sequence, but does not do it efficiently enough to stop the fire spread. For instance, if the agent prioritizes exploration to find the person, then the fire might spread until beyond a critical point where it cannot be extinguished within the horizon.
  • Catastrophic Failure Misunderstanding: Even after the LM has completed the task, it can misunderstand a later failure as a cause for continuing the task. For instance, we have observed that if the LM has successfully dropped off a person, but then fails to pick them up again (which is expected since they’re already dropped), it will think the task is incomplete.
Appendix H Baselines
@@
While there are a lot of impressive LLM-based multi-agent planners as mentioned in Table 1, they vary in the assumptions about access to information about the environment. We were not able to find the official codebase for the Safe Multi-Agent Planning with Conformal Prediction Wang et al. (2024) and TwoStep Singh et al. (2024). We describe the prompts used for our model as well as every baseline. Note that we show the prompt for the 2-agent case, but it is easily modified to generalize to the $n$-agent case. The italics and bolding added for emphasis.
H.1 LLaMAR
@@
description of what failed,
Bob’s observation: list of objects the Bob is observing,
@@ H.6 CoELA
Appendix I Open Source VLMs
@@
  • Idefics 2 Laurençon et al. (2023, 2024): We use the 8B base model fine-tuned on a mixture of supervised and instruction datasets (text-only and multimodal datasets) from HuggingFace. The weights were downloaded from https://huggingface.co/HuggingFaceM4/idefics2-8b with the commit #2c031da2dc71f3ac989f9efa9b8ff476df3842c0. We chose Idefics because it is able to take multiple images as input similar to GPT-4V and reason on them.
  • LLaVA Liu et al. (2023b): We use the 7B model t trained by fine-tuning LLaMA/Vicuna on GPT-generated multimodal instruction-following data. The weights were downloaded from https://huggingface.co/llava-hf/llava-1.5-7b-hf with the commit # 05ae2434cbb430be33edcba0c5203e7023f785b7.
  • CogVLM Hong et al. (2023a): We use the 18B model. The weights were downloaded from https://huggingface.co/THUDM/cogagent-chat-hf with the commit # d519da3b191401234f4bd86ce1c287c61bc276a3.
Appendix J SentenceBERT fine-tuning